OpenStudy (anonymous):

Segment AB has point A located at (4, 2). If the distance from A to B is 3 units, which of the following is the coordinate for point B?

OpenStudy (anonymous):

(5, 2) (1, 5) (4, -1) (-2, 2)

OpenStudy (anonymous):

@paki Can you help?

OpenStudy (anonymous):

@hanner_B_nanner @Orion1213

OpenStudy (anonymous):

@camerondoherty

OpenStudy (anonymous):

distance formula will do...

OpenStudy (anonymous):

idk how to apply it

OpenStudy (anonymous):

\[d=\sqrt{(x_2-x_1)^2+(y_2-2_1)^2}\]

OpenStudy (anonymous):

d=3 units and point A (4,2) is given...

OpenStudy (anonymous):

is it -2,2

OpenStudy (anonymous):

we can check by plugging in the values...

OpenStudy (anonymous):

\[d=\sqrt{(4-(-2))^2+(2-2)^2}=\sqrt{(6)^2+0}=\sqrt{36}=6~units\]therefore it is not the correct answer....

OpenStudy (anonymous):

... wait i'm recalling another formula in solving this problem...

OpenStudy (anonymous):

oh ok...

OpenStudy (anonymous):

1,5 maybe

OpenStudy (anonymous):

it is 4,1... 'cause \(d=\sqrt{(4-4)^2+(2-(-1))^2}=\sqrt{0+9}=3\) units...

OpenStudy (anonymous):

i'm sorry, it should be (4,-1)...

OpenStudy (anonymous):

thanks :) can you help w another ? @Orion1213

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at C (−2, 4), D (−3, 1), E (1, 0), and F (−1, 2). Round your answer to the nearest hundredth when necessary. 10 units 10.77 units 11.52 units 12.35 units

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

10.77?

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

\[d_{CD}=\sqrt{(-3-(-2))^2+(1-4)^2}=\sqrt{1+9}=\sqrt{10}\] \[d_{DE}=\sqrt{(1-3)^2+(-2-(-1))^2}=\sqrt{4+1}=\sqrt{5}\] \[d_{EF}=\sqrt{(-1-1)^2+(2-0)^2}=\sqrt{4+4}=\sqrt{8}\] \[d_{FC}=\sqrt{(-2-(-1))^2+(4-2)^2}=\sqrt{1+4}=\sqrt{5}\] \[P_{CDEF}=\sqrt{10}+\sqrt{5}+\sqrt{8}+\sqrt{5}=10.46~units\] think you're right... maybe it's my round off....